Compute the discriminate. Then determine the number and type of solutions for the given equation
9 x² -7x+5 =0

Answer by ankor@dixie-net.com(22740)

Compute the discriminate.
Then determine the number and type of solutions for the given equation
9x² - 7x + 5 = 0
:
the equation form: a^2 + b^2 + c
:
the discriminant
D = b^2 - 4*a*c
:
in this equation: a=9; b=-7; c=5
:
D = -7^2 - 4 * 9 * 5
D = 49 - 180
D = -131
:
The rules of the discriminant
D > 0; two distinct real roots
D = 0; one real root (a double root)
D < 0; two imaginary roots, one the conjugate of the other
:
Remember these rules and you will have no trouble with the discriminant
:
You should be able figure out the roots in the given equation now